# Geometric Deep Learning for Computer-Aided Design: A Survey

### v1

@@ Abstract
Geometric Deep Learning techniques have become a transformative force in the field of Computer-Aided Design (CAD), and have the potential to revolutionize how designers and engineers approach and enhance the design process. By harnessing the power of machine learning-based methods, CAD designers can optimize their workflows, save time and effort while making better informed decisions, and create designs that are both innovative and practical. The ability to process the CAD designs represented by geometric data and to analyze their encoded features enables the identification of similarities among diverse CAD models, the proposition of alternative designs and enhancements, and even the generation of novel design alternatives. This survey offers a comprehensive overview of learning-based methods in computer-aided design across various categories, including similarity analysis and retrieval, 2D and 3D CAD model synthesis, and CAD generation from point clouds. Additionally, it provides a complete list of benchmark datasets and their characteristics, along with open-source codes that have propelled research in this domain. The final discussion delves into the challenges prevalent in this field, followed by potential future research directions in this rapidly evolving field.
@@
geometric deep learning, machine learning, graph neural networks, computer aided design, automated CAD design, 2D-3D shape modeling
@@ I Introduction
Deep Learning has revolutionized data processing across different domains spanning from grid-structured data in Euclidean spaces, such as image, video, text, to graph-structured data with complex relationships. Geometric Deep Learning (GDL) encompasses a wide array of neural network architectures ranging from Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), Recurrent Neural Networks (RNNs), to Transformer Networks, all of which encode a geometric understanding of the data such as symmetry and invariance as an inductive bias in their learning process [1]. Over the past decade, GDL methods, including CNNs, GNNs, RNNs, and Transformer Networks, have made remarkable strides in diverse tasks coming from different applications, including computer vision, natural language processing, computer graphics, and bioinformatics. However, the application of GDL methods on complex parametric Computer-Aided Design (CAD) data is rarely studied. Boundary Representation (B-Rep), which is a fundamental data format for CAD models encoding a high level of parametric details in CAD models, can be used to learn the intricate geometric features of such models. Although GDL methods have seen considerable success in analyzing 3D shapes in mesh [2, 3], voxel [4, 5, 6], and point cloud [7, 8] data formats, extracting feature representations directly from B-Rep data poses challenges. Converting B-Rep data to conventional formats like triangle meshes is not only computationally expensive but also leads to information loss [9]. Hence, learning feature representations directly from B-Rep data becomes imperative, ensuring an efficient capture of the most representative geometric features in CAD models for subsequent analyses without the burden of extensive computation and memory usage.
@@ II-A2 Terminology in CAD Design
Model
2D or 3D representation of a real-world object or system created within a CAD software.
B-Rep
Boundary Representation, a data format for geometrically describing objects by representing their topological components such as surfaces, edges, vertices, and the relationship between them.
Sketch
2D drawing of an object, served as the basis for creating 3D models, made of lines, curves, and other basic geometric shapes.
Extrude
A CAD operation for expanding a 2D drawing along a specific dimension to create a 3D model.
Primitive
A basic 2D/3D geometric shape which serves as backbone for more complex designs. For 2D sketches, primitives are points, lines, arcs, circles, ellipses and polygons. For 3D shapes primitives are cubes, spheres, cylinders, cones.
Constraint
Geometric relationship, such as coincident, tangent, perpendicular, between primitives in a 2D/3D design.
Point
Basic geometric entity representing a specific location in space with 3D coordinates X, Y, Z.
Line
A straight path between two points in a 3D space.
Circle
A closed curve defined by its center and radius.
Arc
A curved line, as a part of a circle or ellipse, defined by a center, start and end points.
Loop
A closed sequence of curves forming the boundary of a surface.
Profile
A closed loop made by a collection of curves joint together.
Face
A 2D surface bounded by a profile.
Edge
A piece of a curve bounded by two points, defining the boundaries of a face.
Shell
A collection of connected faces with joint boundaries and points.
Body/Part
A 3D shape created by extruding a face.
Solid
A 3D shape that is used as a building block to design an object.
Component
Building blocks of assemblies containing one or more bodies (parts).
Assembly
A collection of connected bodies (or parts) to create a larger object.
Joint
The connection, defined by coordinate and axis, between bodies (parts) to make an assembly.
Topology
The structure of points, curves, faces and other geometric components in a 3D model.
Wireframe
Visual representation of 3D B-Rep models, showing the structure of the object with only lines and curves. Also known as skeletal representation of 3D objects.
Rendering
The process of visualizing CAD models to simulate their real-world appearance.
Surface Normal
A unit vector perpendicular to a surface at a specific point.
@@
STEP
Original parametric B-Rep format containing explicit description of the topology and geometry information of the CAD models.
.txt, .step, .smt
STL
Standard Tessellation Language, a format describing 3D surfaces of an object with triangular facets (meshes).
.stl
Obj
A format describing 3D surface geometry using a triangular meshes.
.obj
Statistics
Statistical infromation of CAD models.
.yml, .txt
Features
A description of the properties of surfaces and curves with references to the corresponding vertices and faces of the discrete triangle mesh representation.
.yml
Image
Visualization of objects produced by rendering process.
.png
PCD
Point Cloud Data, a collection of points in the 3D space representing the structure of a 3D model.
.ASC, .PTX, .OBJ
@@ V CAD Construction with Generative Deep Learning
CAD construction with generative deep learning involves leveraging advanced GDL methods to automatically generate or assist in the creation of parametric CAD models. These approaches can support designers in various ways to streamline the design process. This includes tasks like generating or auto-completing sketches, as well as generating CAD operations to construct a 3D model based on the designed sketch. In this section, we categorize these methods into five distinct groups: 1) methods focusing on 2D sketches, aiming to automate the sketching process in a 2D space as an initial step before developing 3D CAD models, 2) methods targeting 3D CAD model reconstruction given the sketch of the model, 3) methods generating CAD construction sequences, specifically focused on sketch and extrude operations, for CAD construction, 4) methods which perform direct B-Rep synthesis to generate 3D CAD models, and 5) methods generating 3D CAD models from Point Cloud data.
@@ V-C1 DeepCAD [17]
Figure 13: Schematic representation of the DeepCAD model architecture [17]. The model, depicted on the left, is a transformer autoencoder network trained in an unsupervised manner. It takes a sequence of CAD commands as input and reconstructs them. The detailed structure of this autoencoder is shown on the right, where $\left\{C_{1},C_{2},...,C_{N}\right\}$ tokens represent the input CAD command sequence, and $\left\{C_{1},C_{2},...,C_{N}\right\}$ are the reconstructed commands.
@@ V-D2 CADParser [16]
As highlighted in the overview of existing 3D CAD generative models, the reverse engineering of CAD shapes has primarily been explored through methods utilizing either CAD sketches, B-Rep data or sequences of CAD commands as input. These approaches have made promising advances so far in reconstructing and generating CAD models. However, an equally crucial aspect is generating CAD data from alternative raw geometric data modalities, such as point clouds. This consideration also aligns with the future work outlined by the DeepCAD method. The ability to seamlessly convert diverse data representations into CAD models opens new avenues for real-world applications, bridging the gap between different data modalities and expanding the utility of CAD technology. This becomes especially critical in scenarios where new variations of a physical object are needed or when repairing a machinery object without access to the corresponding CAD model. This can be particularly challenging in instances where the object predates the digital era of manufacturing. In these scenarios, the process typically begins with scanning the object using a 3D sensor, which generates a point cloud. Subsequently, the acquired point cloud data needs to be decomposed into a collection of geometric primitives, such as curves or surfaces, to be interpreted by CAD software. The traditional three-step procedure [114] involves converting the point cloud into a mesh, explaining it through parametric surfaces to create a solid (B-Rep), and inferring a CAD program. Recent advancements in fitting primitives to point clouds [115, 116, 117] have managed to bypass the initial step of converting point clouds to meshes. Nevertheless, a notable limitation of these methods is their reliance on a finite set of fixed and disjoint primitives, which poses a challenge for convenient shape editing in subsequent steps. Addressing this, the recently proposed Point2Cyl [32] method frames the problem as an Extrusion Cylinder decomposition task, leveraging a neural network to predict per-point extrusion instances, surface normals, and base/barrel membership. These geometric proxies can then be used to estimate the extrusion parameters through differentiable and closed-form formulations. In this method, an Extrusion Cylinder is considered a fundamental primitive that signifies an extruded 2D sketch, characterized by parameters such as the extrusion axis, center, sketch, and sketch scale, which can be used to represent 3D CAD models. The terms base and barrel are utilized to denote specific surfaces of an extrusion cylinder, representing the base/top plane and the side surface of the extrusion cylinder, respectively. Figure 14 shows an example of Extrusion Cylinder which is obtained by extruding a circle as the base 2D sketch.
Figure 14: An example of Extrusion Cylinder. The 2D sketch is a circle here (right) which is extruded to make the Extrusion Cylinder (left) [32].

### v2

@@ Abstract
Geometric Deep Learning techniques have become a transformative force in the field of Computer-Aided Design (CAD), and have the potential to revolutionize how designers and engineers approach and enhance the design process. By harnessing the power of machine learning-based methods, CAD designers can optimize their workflows, save time and effort while making better informed decisions, and create designs that are both innovative and practical. The ability to process the CAD designs represented by geometric data and to analyze their encoded features enables the identification of similarities among diverse CAD models, the proposition of alternative designs and enhancements, and even the generation of novel design alternatives. This survey offers a comprehensive overview of learning-based methods in computer-aided design across various categories, including similarity analysis and retrieval, 2D and 3D CAD model synthesis, and CAD generation from point clouds, and single/multi-view images. Additionally, it provides a complete list of benchmark datasets and their characteristics, along with open-source codes that have propelled research in this domain. The final discussion delves into the challenges prevalent in this field, followed by potential future research directions in this rapidly evolving field.
@@
Computer-aided design, geometric deep learning, graph neural networks, machine learning
@@ I Introduction
Deep Learning has revolutionized data processing across different domains spanning from grid-structured data in Euclidean spaces, such as image, video, text, to graph-structured data with complex relationships. Geometric Deep Learning (GDL) encompasses a wide array of neural network architectures ranging from Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), Recurrent Neural Networks (RNNs), to Transformer Networks, all of which encode a geometric understanding of the data such as symmetry and invariance as an inductive bias in their learning process [1]. Over the past decade, GDL methods, including CNNs, GNNs, RNNs, and Transformer Networks, have made remarkable strides in diverse tasks coming from different applications, including computer vision, natural language processing, computer graphics, and bioinformatics. However, the application of GDL methods on complex parametric Computer-Aided Design (CAD) data is rarely studied. Boundary Representation (B-Rep), which is a fundamental data format for CAD models encoding a high level of parametric details in CAD models, can be used to learn the intricate geometric features of such models.
Although GDL methods have seen considerable success in analyzing 3D shapes in mesh [2, 3], voxel [4, 5, 6], and point cloud [7, 8] data formats, extracting feature representations directly from B-Rep data poses challenges. Converting B-Rep data to conventional formats like triangle meshes is not only computationally expensive but also leads to information loss [9]. Hence, learning feature representations directly from B-Rep data becomes imperative, ensuring an efficient capture of the most representative geometric features in CAD models for subsequent analyses without the burden of extensive computation and memory usage.
@@ II-A2 Terminology in CAD Design
Model 2D or 3D representation of a real-world object or system created within a CAD software.
B-Rep Boundary Representation, a data format for geometrically describing objects by representing their topological components such as surfaces, edges, vertices, and the relationship between them.
Sketch 2D drawing of an object, served as the basis for creating 3D models, made of lines, curves, and other basic geometric shapes.
Extrude A CAD operation for expanding a 2D drawing along a specific dimension to create a 3D model.
Primitive A basic 2D/3D geometric shape which serves as backbone for more complex designs. For 2D sketches, primitives are points, lines, arcs, circles, ellipses and polygons. For 3D shapes primitives are cubes, spheres, cylinders, cones.
Constraint Geometric relationship, such as coincident, tangent, perpendicular, between primitives in a 2D/3D design.
Point Basic geometric entity representing a specific location in space with 3D coordinates X, Y, Z.
Line A straight path between two points in a 3D space.
Circle A closed curve defined by its center and radius.
Arc A curved line, as a part of a circle or ellipse, defined by a center, start and end points.
Loop A closed sequence of curves forming the boundary of a surface.
Profile A closed loop made by a collection of curves joint together.
Face A 2D surface bounded by a profile.
Edge A piece of a curve bounded by two points, defining the boundaries of a face.
Shell A collection of connected faces with joint boundaries and points.
Body/Part A 3D shape created by extruding a face.
Solid A 3D shape that is used as a building block to design an object.
Component Building blocks of assemblies containing one or more bodies (parts).
Assembly A collection of connected bodies (or parts) to create a larger object.
Joint The connection, defined by coordinate and axis, between bodies (parts) to make an assembly.
Topology The structure of points, curves, faces and other geometric components in a 3D model.
Wireframe Visual representation of 3D B-Rep models, showing the structure of the object with only lines and curves. Also known as skeletal representation of 3D objects.
Rendering The process of visualizing CAD models to simulate their real-world appearance.
Surface Normal A unit vector perpendicular to a surface at a specific point.
@@
STEP Original parametric B-Rep format containing explicit description of the topology and geometry information of the CAD models. .txt, .step, .smt
STL Standard Tessellation Language, a format describing 3D surfaces of an object with triangular facets (meshes). .stl
Obj A format describing 3D surface geometry using a triangular meshes. .obj
Statistics Statistical infromation of CAD models. .yml, .txt
Features A description of the properties of surfaces and curves with references to the corresponding vertices and faces of the discrete triangle mesh representation. .yml
Image Visualization of objects produced by rendering process. .png
PCD Point Cloud Data, a collection of points in the 3D space representing the structure of a 3D model. .ASC, .PTX, .OBJ
@@ II-B Geometric Deep Learning (GDL)
The field of geometric deep learning for CAD has progressed through distinct stages, each addressing limitations of the previous generation of models in handling the structured, parametric, and semantic-rich nature of CAD data. Early approaches such as MeshCNN [54] operated on triangle meshes using edge-based convolutions to capture local surface structures. While effective for organic shapes with well-formed manifold meshes, these models struggle with the topological complexity and degeneracies typical in industrial CAD exports, such as T-vertices, non-manifold edges, and inconsistent triangulations.
To improve robustness, point-based models like PointNet [8] and DGCNN [55] were introduced, which treat 3D shapes as unordered sets of points and learn global or local geometric features without relying on mesh connectivity. These methods offer greater resilience to topological inconsistencies and mesh noise. However, they come with a critical drawback: they discard the explicit topological and parametric structure of CAD models, including analytic surface types, like planes, cylinders, design features, and constraint relationships. As such, while useful for tasks like classification or segmentation, they fall short in applications that require understanding or manipulation of design intent.
To bridge this gap, a new class of graph-based models emerged that leverage the native boundary representation (B-Rep) topology inherent to CAD data. These Graph Neural Networks (GNNs) operate on structured graphs built from B-Rep entities, faces, edges, and vertices, enabling explicit encoding of parametric surfaces, topological constraints, and hierarchical relationships between geometric features. Models such as UV-Net [9], DeepCAD [17] and BRepNet [14] exemplify this direction, offering better adaptability to CAD workflows and supporting tasks like feature recognition, constraint extraction, and semantic segmentation on structured geometry.
Building upon this foundation, recent Transformer-based models such as SketchGen [26], CAD As Language [27], PolyGen [56] and other recent generative methods reviewed in this paper as illustrated in 1 represent the current state of the art. These models operate directly on tokenized B-Rep elements or design operation sequences, capturing long-range dependencies and supporting learning across rich semantic and geometric contexts. They are not only capable of encoding exact geometry, but also learn design history, manufacturability constraints, and support editability and reverse engineering, aligning closely with how CAD software encodes and manipulates geometry.
This technological evolution, from mesh- and point cloud-based methods to graph- and transformer-based architectures, demonstrates a clear trajectory toward CAD-native learning, with increasing alignment to the structural, semantic, and engineering-specific properties of design data.
II-C Graph neural networks (GNNs)
@@ V CAD Construction with Generative Deep Learning
CAD construction with generative deep learning involves leveraging advanced GDL methods to automatically generate or assist in the creation of parametric CAD models. These approaches can support designers in various ways to streamline the design process. This includes tasks like generating or auto-completing sketches, as well as generating CAD operations to construct a 3D model based on the designed sketch. In this section, we categorize these methods into six distinct groups: 1) methods focusing on 2D sketches, aiming to automate the sketching process in a 2D space as an initial step before developing 3D CAD models, 2) methods targeting 3D CAD model reconstruction given the sketch of the model, 3) methods generating CAD construction sequences, specifically focused on sketch and extrude operations, for CAD construction, 4) methods which perform direct B-Rep synthesis to generate 3D CAD models, 5) methods generating 3D CAD models from Point Cloud data, and 6) methods generating CAD models from images.
@@ V-C1 DeepCAD [17]
Figure 13: Schematic representation of the DeepCAD model architecture [17]. The model, is a transformer autoencoder network trained in an unsupervised manner. It takes a sequence of CAD commands $\left\{C_{1},C_{2},...,C_{N}\right\}$ as input and generates the reconstructed commands $\left\{\hat{C}_{1},\hat{C}_{2},...,\hat{C}_{N}\right\}$ as output.
@@ V-C2 SkexGen [31]
V-C3 Diffusion-CAD [32]
Diffusion-CAD [32] introduces a diffusion-based generative model for synthesizing 3D CAD construction sequences with a particular emphasis on fine-grained user control and structural validity. In contrast to prior methods such as DeepCAD, and SkexGen, which either lack explicit controllability or rely on discrete sampling from latent codebooks, Diffusion-CAD formulates CAD command generation as a continuous denoising diffusion process over embedded command sequences. This approach enables greater control and a smoother sampling landscape for generation.
In Diffusion-CAD, each CAD model is represented as a fixed-length sequence of $N=60$ commands, where each command consists of a type identifier (e.g., line, arc, circle, extrude) and a set of up to 17 geometric parameters. The commands are embedded into continuous vectors and distorted with Gaussian noise across time steps in the forward diffusion process. A Transformer-based denoising network is then trained to recover the original command vectors, which are finally projected back into discrete command tokens. This framework allows for generating valid and diverse CAD models from noise, while also supporting several control mechanisms during generation. A key contribution of Diffusion-CAD is its support for multiple forms of user-defined control, including:
  • Command-type control: constraining the generation to specific types of primitives (e.g., only circles and extrusions),
  • Dimension control: fixing specific geometric parameters such as radii or angles,
  • Partial sketch completion: completing full CAD sequences given a user-provided prefix,
  • Structural control: enforcing relationships like symmetry, perpendicularity, or collinearity via constraint-satisfaction modules,
  • Class-conditional generation: generating shapes from specific semantic classes using classifier-guided diffusion.
Compared to DeepCAD [17], which struggles with structural consistency, and SkexGen [31], which lacks fine control over geometry, Diffusion-CAD demonstrates more robust and flexible generation. It marks a significant step forward in CAD sequence generation, combining the strengths of deep generative modeling with explicit, user-guided controllability to produce valid and expressive CAD models.
@@ V-D2 CADParser [16]
V-D3 BrepGen [33]
BrepGen [33] introduces a diffusion-based generative model that directly synthesizes complete B-Rep CAD models, addressing the limitations of prior methods such as SolidGen, which often lacked geometric expressiveness and struggled with topological validity. BrepGen departs from the autoregressive modeling used in SolidGen and instead employs a structured latent geometry representation, where the entire B-Rep is organized as a hierarchical tree: solids contain faces, faces contain edges, and edges contain vertices. Each node in this tree includes both topological and geometric features, which are encoded via latent embeddings.
One of the key innovations in BrepGen is its approach to handling topology. Rather than explicitly generating adjacency matrices, it duplicates topological elements across branches of the tree and reconstructs watertight solids by merging equivalent nodes. This implicit treatment simplifies the generative process but relies on post-hoc merging, which may occasionally lead to inconsistencies or self-intersections. On the geometric side, BrepGen supports curved and freeform surfaces by encoding geometry through UV-sampled surface and curve point clouds. These are compressed via variational autoencoders (VAEs), allowing for a more compact representation of complex geometry than SolidGen, which is limited to simpler surfaces like planes and cylinders. Each part of the geometry (vertices, edges, and faces) is generated through separate denoising diffusion models, enabling expressive and controllable shape synthesis.
BrepGen demonstrates high performance on DeepCAD and ABC datasets, outperforming both DeepCAD’s sketch-based approach and SolidGen in terms of geometric integrity (e.g., Chamfer distance), uniqueness, novelty, and topological validity. It also supports conditional generation tasks such as shape interpolation and class-guided synthesis. However, the reliance on latent geometry codes and implicit topology inference introduces certain limitations, such as reduced interpretability and dependency on successful merging of duplicated elements. These limitations are addressed by subsequent method, DTGBrepGen [34], which explicitly decouple topology and geometry generation to ensure structural consistency and improve control over the synthesis process.
V-D4 DTGBrepGen [34]
DTGBrepGen [34] builds upon the strengths of BrepGen [33] by addressing its key limitations, specifically, the implicit and post-hoc treatment of B-Rep topology. Instead of jointly modeling geometry and topology within a latent tree structure, DTGBrepGen introduces a principled decoupling of topology and geometry into two separate stages, resulting in higher structural validity and more accurate geometric outputs.
The first stage of DTGBrepGen focuses entirely on topology generation, constructing valid edge-face and edge-vertex adjacency matrices using two Transformer-based encoder-decoder networks. This step ensures compliance with essential topological constraints, such as each edge connecting two distinct vertices and forming closed loops within each face. Unlike BrepGen, which infers these relationships indirectly through merging duplicated nodes, DTGBrepGen generates them explicitly and in a controlled manner. Once a valid topological structure is generated, the second stage of the pipeline uses Transformer-based diffusion models to sequentially generate geometric attributes: vertex coordinates, edge curves, and face surfaces. Geometries are represented using cubic and bicubic B-splines, enabling mathematically precise modeling of complex curves and surfaces. This choice avoids the need for surface sampling and point cloud-based encoders used in BrepGen, improving both integrity and interpretability.
DTGBrepGen achieves superior performance across multiple benchmarks, on datasets such as DeepCAD and ABC. It also supports class-conditional and point cloud-conditioned generation, showcasing its generalizability across design tasks. By generating valid topologies first and conditioning geometry on them, DTGBrepGen produces watertight, structurally consistent CAD models with more accurate and editable geometries. This modular design also enhances interpretability and lays the groundwork for future extensions involving hierarchical modeling or integration with semantic constraints.
@@ V-D4 DTGBrepGen [34]
As highlighted in the overview of existing 3D CAD generative models, the reverse engineering of CAD shapes has primarily been explored through methods utilizing either CAD sketches, B-Rep data or sequences of CAD commands as input. These approaches have made promising advances so far in reconstructing and generating CAD models. However, an equally crucial aspect is generating CAD data from alternative raw geometric data modalities, such as point clouds. This consideration also aligns with the future work outlined by the DeepCAD method. The ability to seamlessly convert diverse data representations into CAD models opens new avenues for real-world applications, bridging the gap between different data modalities and expanding the utility of CAD technology. This becomes especially critical in scenarios where new variations of a physical object are needed or when repairing a machinery object without access to the corresponding CAD model. This can be particularly challenging in instances where the object predates the digital era of manufacturing. In these scenarios, the process typically begins with scanning the object using a 3D sensor, which generates a point cloud. Subsequently, the acquired point cloud data needs to be decomposed into a collection of geometric primitives, such as curves or surfaces, to be interpreted by CAD software. The traditional three-step procedure [121] involves converting the point cloud into a mesh, explaining it through parametric surfaces to create a solid (B-Rep), and inferring a CAD program. Recent advancements in fitting primitives to point clouds [122, 123, 124] have managed to bypass the initial step of converting point clouds to meshes. Nevertheless, a notable limitation of these methods is their reliance on a finite set of fixed and disjoint primitives, which poses a challenge for convenient shape editing in subsequent steps. Addressing this, the recently proposed Point2Cyl [35] method frames the problem as an Extrusion Cylinder decomposition task, leveraging a neural network to predict per-point extrusion instances, surface normals, and base/barrel membership. These geometric proxies can then be used to estimate the extrusion parameters through differentiable and closed-form formulations. In this method, an Extrusion Cylinder is considered a fundamental primitive that signifies an extruded 2D sketch, characterized by parameters such as the extrusion axis, center, sketch, and sketch scale, which can be used to represent 3D CAD models. The terms base and barrel are utilized to denote specific surfaces of an extrusion cylinder, representing the base/top plane and the side surface of the extrusion cylinder, respectively.
@@
Recently, Ma et al. [38] proposed a novel method for reconstructing CAD programs from point clouds using a multimodal diffusion model. Unlike previous methods focused on geometric primitives or B-Rep structures, this approach tokenizes CAD operations and sketches, enabling the generation of editable CAD construction sequences. Through a volume-based diffusion process, it progressively reconstructs human-like design steps, bridging raw 3D data and structured CAD workflows. This method complements prior works by elevating the abstraction from geometry to procedural modeling, making it particularly suitable for interactive reverse engineering tasks.
V-F 3D CAD Generation from Image
While CAD generative methods discussed in the previous four categories rely on structured inputs such as sketches, point clouds, or construction sequences, recent advancements have explored generating CAD models directly from unstructured RGB images. These methods attempt to infer structured, editable CAD representations, such as B-reps, from single or multi-view images. This new direction bridges the gap between 2D visual understanding and 3D modeling, enabling applications in reverse engineering, industrial inspection, and conceptual design from photos.
CADDreamer [39] proposes a pipeline that reconstructs a structured, watertight B-rep CAD model from a single-view RGB image. Unlike traditional 3D reconstruction pipelines that produce meshes or voxels, CADDreamer directly targets CAD-suitable representations by combining primitive-aware diffusion models with geometric optimization. The core idea of this method is to first generate normal maps and semantic primitive segmentation maps (covering planes, cylinders, cones, spheres, and feature curves) using a multi-view diffusion model. These are then fused across virtual views to build a multi-surface representation of the object. A differentiable renderer then estimates a coarse mesh from the normal maps, while the primitive maps guide segmentation and graph-cut-based patching. The system then performs primitive fitting and topological trimming to reconstruct a valid and editable B-rep with sharp features and semantic structure. The main advantage of this method is its ability to operate on single-view RGB input, eliminating the need for explicit geometry sensing, while producing compact and editable B-rep models. However, its reconstruction quality is dependent on accurate segmentation and primitive fitting, and the method currently lacks support for parametric design history and user-guided editing.
CADCrafter [40] addresses the same image-to-CAD problem, but instead of generating geometric models directly, it focuses on recovering CAD construction sequences (e.g., sketches and extrusions) that can be imported into CAD software such as Fusion360. This approach reflects real-world design workflows and enables fully editable, parametric outputs. This method uses a geometry-conditioned latent diffusion transformer that generates a sequence of CAD tokens, which are then decoded into construction commands. The conditioning is achieved by a geometry encoder trained to be texture-invariant, using estimated depth and normal maps derived from the input image. This helps bridge the domain gap between synthetic CAD renders (used during training) and real-world, unconstrained images. To improve correctness, CADCrafter integrates a code-checking module that compiles and validates the generated CAD commands. Furthermore, it introduces a new RealCAD dataset, which contains real-world images paired with command sequences, enabling better generalization. The main advantage of CADCrafter is its ability to operate on real-world, unconstrained images, supporting both single- and multi-view inputs. It demonstrates strong performance under domain shift, thanks to its geometry-based conditioning, and the use of preference optimization. However, the method is currently limited to sketch-and-extrude workflows and does not yet support more complex or advanced CAD operations.
CADDreamer and CADCrafter demonstrate how deep generative models can automate parametric modeling tasks by interpreting visual content, producing CAD outputs that are both human-readable and machine-executable.